You and your friend live 12 miles apart. you leave home at the same time and travel toward each other. you walk at a rate of 4 miles per hour and your friend bicycles at a rate of 11 miles per hour.
a. How long after you leave home will you meet? how far from home will each of you be?
b. Suppose your friend bicycles at a rate of 12 miles per hour. how much sooner will you meet? How far from home will each of you be?

Answer by mananth(16949)

Distance = 12
speed = 4+11 = 15 mph ( since you both are travelling towards each other)
time = 12/15 * 60 = 48 minutes ===> 4/5 hours
In 4/5 hours you wouls have traveled 4/5 * 4 = 16/5 miles
friend would have traveled = 4/5 * 11 = 44/5 miles
..
I think you can do the next part